OpenStudy (anonymous):

Find the roots of the quadratic equation x2 – 8x = 9 by completing the square. Show your work.

OpenStudy (anonymous):

\[x^2-8x=9\] \[x^2-8x+(8/2)^2=9(8/2)^2\] \[x^2-8x+16=25\] \[\sqrt{(x-4)^2}=\sqrt{25}\] \[x-4=\pm5\] x=1,x=-9

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

welcome :-)

OpenStudy (anonymous):

very helpfull